OpenStudy (anonymous):

What is the axis of symmetry for the function? y = x2 + 3x - 4 A. y = -2 B. y = -1.5 C. x = -2 D. x = -1.5

OpenStudy (anonymous):

y'=2x+3=0, or x=-3/2, or D.

OpenStudy (anonymous):

If your equation is in standard form, then the formula for the axis of symmetry is: x = -b/2a from the general standard form equation y = ax2+bx + c http://www.mathwarehouse.com/geometry/parabola/axis-of-symmetry.php

OpenStudy (anonymous):

Look at you, being helpful and all.

OpenStudy (anonymous):

THanks. I try my best to work in a team.